# Learning When to Plan: Efficiently Allocating Test-Time Compute for LLM Agents

### v1

@@ 1 Introduction
Figure 1: Dynamic planning strategies across environments and training stages. (a-b) Zero-shot results showing optimal "Goldilocks" planning frequency in Crafter and POGS (100 seeds, bars=standard-error). (c-d) SFT results demonstrating planning agents’ improved performance with lower KL divergence from base model. (e-f) RL results where SFT-primed planning agents are more sample efficient than non-planning baselines and more consistently reach complex achievements.
@@
    Each task has a "Goldilocks" frequency for planning that clearly outperforms naive strategies of always planning or never planning.
@@
Our work provides clear evidence that dynamic planning facilitates effective allocation of test-time compute in sequential decision making environments, showing that LLM agents can be trained to use additional computational resources intelligently. The ability to steer such planning agents—now capable enough to complete Crafter by collecting diamond under human guidance—marks a significant step towards safer and more collaborative LLM agents. Together, these findings suggest a promising path towards more capable, efficient, interpretable, and steerable agentic systems.
@@ 2 Related Work
Historically, much progress in sequential decision making has involved systems that explicitly look ahead before acting. Monte Carlo Tree Search (MCTS)(Coulom, 2006), combined with deep neural networks, has driven landmark systems like AlphaGo and MuZero(Silver et al., 2017b, a; Schrittwieser et al., 2020). Model Predictive Control (MPC) iteratively plans short horizons, adapting to new observations (Mayne et al., 2000). Similarly, model-based RL methods, such as World Models, PlaNet, and Dreamer, use imagination rollouts in latent space for effective planning and learning (Ha and Schmidhuber, 2018; Hafner et al., 2019, 2020). Collectively, these approaches underscore the strength of explicit planning, particularly when accurate internal or environmental models are available.
@@
Large language models have demonstrated significant reasoning capabilities, particularly through techniques like chain-of-thought (CoT) prompting (Wei et al., 2022). ReAct extends CoT into sequential settings, explicitly prompting models to reason before acting (Yao et al., 2023b). Similar reasoning methods include self-reflective prompting (Shinn et al., 2023; Wang et al., 2023b; Hao et al., 2023; Besta et al., 2024; Yao et al., 2023a), automated prompt tuning (Fernando et al., 2024; Hu et al., 2025), and strategic planning demonstrated by CICERO in the Diplomacy game († et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR). However, frequent replanning can cause behavioural instability, analogous to RL frame-skipping strategies that advocate less frequent action repetition for improved exploration and consistency (Sharma et al., 2017; Kalyanakrishnan et al., 2021). Recent studies also show diminishing returns and increased brittleness from excessive reasoning (Stechly et al., 2024; Mizrahi et al., 2024; Sui et al., 2025), emphasizing the need for adaptive planning mechanisms. Behaviour cloning with LLMs on data that includes textual reasoning between actions has been shown to help imitation learning (Yang et al., 2022; Hu and Clune, 2023).
@@
Connection to Hierarchical Reinforcement Learning
Previous HRL work involving LLMs focused on discovering low-level skills (Wang et al., 2023a; Klissarov et al., ) and making high-level decisions about which skills to use (Ahn et al., 2022; Klissarov et al., 2025). Traditionally, HRL frameworks require the high-level controller to be invoked at every fixed interval or even every time step to decide which option or sub-policy to execute next; this repeated querying is common in modular HRL architectures and is a computational bottleneck (Sutton et al., 1999; Bacon et al., 2017). This work diverges from the conventional HRL paradigm by focusing not on what sub-policy to choose, but on when to engage the costly planning process. By training an agent to dynamically decide when to plan, our approach directly addresses the computational constraints of using LLMs for strategic reasoning, optimizing the allocation of test-time compute.
@@ 4.1 Environments
To evaluate dynamic planning across different conditions, we selected two complementary environments. First, Partially Observable Graph Search (POGS) is our custom synthetic environment designed to isolate planning under uncertainty. Agents navigate procedurally generated graph using only local observations, which require adaptive replanning upon discovering new nodes or dead ends. POGS allows measurement of exploration efficiency via backtracking statistics. Second, Crafter (Hafner, 2022) is a complex 2D grid-world, long-horizon benchmark inspired by Minecraft. It demands multi-scale planning for survival, resource management, and crafting, testing both short-term tactical decisions and long-term strategic choices. Interaction in both environments occurs via natural language. Figure 3 showcases the environments, and full technical details are provided in the Appendix B.
@@ 4.2 Evaluation Protocol
We utilize the BALROG benchmark (Paglieri et al., 2025a) for standardized agent evaluation and environment interaction. At each timestep $t$, the agent receives its history and current observation $o_{t}$ within a chat template, guided by a system prompt outlining the task. The agent’s response must include a natural language action command $a_{t}$. Our dynamic planning agents are instructed to decide at each step whether to plan. If they choose not to plan, they output only the action command [Action]. If they choose to plan, they output the plan followed by the action, using the format <plan> [natural language plan] </plan> [Action]. BALROG parses this output, identifying a planning decision ($d_{t}=1$) if the <plan> block is present and using its content as the current plan $p_{t}$ in subsequent context. The [Action] is always extracted and executed. Fallback mechanisms ensure robustness against invalid outputs. Appendix A provides detailed prompts.
@@ 4.4 Supervised Fine-Tuning (SFT)
SFT Priming: The Llama-3.1-8B model is fine-tuned on this data, aligning the SFT process with the target RL configuration. For the SFT+RL plan dynamically agent, SFT targets use the dynamic format – <plan>...</plan> [Action] if the teacher planned at that step, otherwise just [Action] – along with a dynamic prompt encouraging this choice. Conversely, for the SFT+RL no plan agent, SFT targets include only the actions ([Action], with all plan blocks removed), paired with a naive prompt focused solely on action prediction. This prepares the model appropriately for the subsequent RL phase. More details on the prompt in the Appendix A.
@@ 4.5 Reinforcement Learning (RL)
We then use Proximal Policy Optimization (PPO) (Schulman et al., 2017) to fine-tune Llama-3.1-8B-Instruct agents in Crafter, optimizing task rewards possibly adjusted for planning costs (Sec. 3.3). We compare four key configurations:
@@ 5.1 Zero-Shot Evaluation
We evaluated the zero-shot performance of Llama-3.3-70B-Instruct at various planning frequencies, as shown in Figure 1 (a) and (b) for POGS and Crafter. Average task progression is plotted against mean output tokens (log scale, proxying computational cost), comparing agents that plan at different intervals-from never to every step.
Results in both environments reveal a non-monotonic relationship: some planning improves over naive agents, but excessive planning reduces performance. Performance peaks at intermediate frequencies (e.g., every 4 steps in Crafter), then declines as planning becomes more frequent. The always-plan approach, similar to standard ReAct, consistently underperforms compared to less frequent planning, despite higher computational cost. This demonstrates an optimal, task-dependent "Goldilocks" zone for planning frequency. This supports our framework (Section 3), which introduces an instability cost ($C_{noise}$) from frequent planning. While planning reduces uncertainty, replanning too often with noisy or inconsistent plans introduces instability, causing inefficient behaviours like oscillating between subgoals or abandoning viable actions.
Quantitative backtracking analysis in POGS (Appendix B) provides clear evidence: the always-plan agent showed the most backtracking, highlighting its tendency to oscillate or revisit states rather than explore efficiently. Agents planning at intermediate frequencies achieved better progression with significantly less backtracking, reflecting more systematic exploration. Crafter shows similar patterns: frequent replanning causes course changes that stall progress and induce inefficiencies.
Our initial attempts to elicit adaptive dynamic planning behaviour directly via complex zero-shot prompting proved challenging and unreliable. Models struggled to consistently interpret instructions like "plan when necessary," often defaulting to fixed patterns. This difficulty underscores the need for learning-based approaches, like SFT and RL, to effectively teach agents this meta-cognitive skill of deciding when to allocate test-time compute for planning.
@@ 5.2 SFT Priming
Having established the importance of planning frequency, we next evaluate our SFT priming stage (Section 4.4), designed to prepare the Llama-3.1-8B model for subsequent RL fine-tuning. We compare the performance of the ‘Primed-Dynamic’ and ‘Primed-Naive’ models during the SFT process itself, evaluating their Crafter progression periodically throughout training (Figure 1c).
The key comparison here is that both models are fine-tuned on datasets derived from the exact same underlying action sequences. The sole difference lies in the target outputs: ‘Primed-Dynamic’ is trained to predict the explicit natural language plans generated by the teacher model (when present) alongside the actions, using the dynamic planning format, while ‘Primed-Naive’ is trained to predict only the actions, with all plan information removed.
Despite the identical action supervision, the ‘Primed-Dynamic’ model consistently achieves higher task progression throughout the SFT phase. This result strongly suggests that access to the explicit natural language plans significantly aids the imitation learning process itself. We hypothesize several contributing factors for this observation:
@@
Explanatory Power: Plans provide semantic context and rationale for subsequent actions, potentially simplifying the behavioural cloning objective by making the action sequence more predictable.
Learning Planning Structure: Exposure to diverse planning data likely helps the model learn generalizable representations related to “look-ahead,” improving performance beyond simple imitation.
@@
Regularization and Grounding: Natural language plans may act as a form of cognitive grounding, constraining the model’s behaviour during fine-tuning and preventing it from deviating too drastically from the base model’s capabilities in unproductive ways.
Supporting the grounding hypothesis, preliminary analyses indicate that the “Primed-Dynamic” model exhibits lower KL divergence from the base pre-trained model during SFT compared to the “Primed-Naive” model (see Figure 1d). This suggests that training with plans may support a less disruptive fine-tuning process. Overall, these findings highlight the value of incorporating explicit reasoning in the form of natural language plans directly into training data – not merely as a target behaviour, but as an effective mechanism to improve the learning process itself.
5.3 RL Fine-Tuning
Following SFT priming, we fine-tuned all agent variants through RL. Our results reveal that SFT-primed agent equipped with dynamic planning (SFT+RL plan dynamically) substantially outperforms the non-planning counterparts (SFT+RL no plan) during the initial phases of gameplay. Notably, qualitative analysis shows that the SFT+RL plan dynamically agent is able to generate and execute plans at multiple levels of abstraction, flexibly shifting between high-level strategic objectives and lower-level tactical actions depending on the demands of the current situation. Furthermore, the agent demonstrates the ability to plan only when necessary, autonomously executing plans and replanning in response to changing circumstances. Qualitative results in Appendix D.
However, as the environment becomes more challenging, the improvement rates for both SFT+RL agents diminish, resulting in eventual plateauing of their progression curves at comparable levels. This plateau suggests that while dynamic planning confers a clear advantage in early and mid-game scenarios, both agent types ultimately encounter similar bottlenecks in late-game survival and high-level achievement. Qualitative analysis attributes this stagnation to persistent deficiencies in survival skills such as acquiring food and water or building a shelter, indicating that further progress likely requires additional training beyond our current computational budget.
When comparing agents initialized from the base Llama-3.1-8B-Instruct model (Base+RL plan dynamically and Base+RL no plan), we observe inferior performance throughout training. Notably, the Base+RL plan dynamically agent exhibits the most limited progression, struggling to leverage planning in the absence of SFT priming. We hypothesize that the 8B model, without prior SFT, may possess insufficient scale or inherent capability to autonomously generate planning traces that are consistently effective for solving the task. Consequently, for models of this scale, structured, expert-like planning examples provided through SFT appear crucial for enabling effective downstream RL optimization and the emergence of dynamic planning behaviour.
Instruction following: Finally, we explore human-agent collaboration by steering agents with human-written plans. The base model exhibits significant grounding issues, often attempting invalid actions that lead to unproductive behaviour and and quick death when fighting a monster. The SFT-primed model, while much more efficient than the base model, still exhibits numerous errors, doesn’t learn from feedback, and struggles to adhere to novel plans, often defaulting to previously learned strategies due to excessive memorization. For example, an SFT agent might ignore a human plan to mine stone, instead prioritizing combat with nearby monsters. In contrast, RL-trained agents both excel at the game and reliably follow external plans; one such agent was successfully guided by human plans to collect diamond and solve Crafter-an achievement unseen in autonomous agents. We also show a complex plan following example in Figure 4, while we refer the reader to the Appendix D for more qualitative results. To quantify the upper bound of performance, we conducted a Best-of-N analysis, comparing the best runs attained by each method across 100 independent runs, and 20 runs for the human collaboration. As shown in Figure 5, the SFT+RL when guided by human plans, successfully solves Crafter by mining a diamond, significantly outperforming all of the methods. This result demonstrates the impact of human steering and further validates the benefits of conditioning the models with plans.
In summary, our RL experiments demonstrate that SFT priming is essential for enabling dynamic planning to emerge in LLM agents. The resulting SFT+RL plan dynamically agent exhibits better sample efficiency and superior performance on advanced metrics. Nevertheless, persistent challenges in late-game survival and the observed performance plateau indicate that further advances will require additional training resources or other innovations. Addressing these limitations will be crucial for developing the next wave of reasoning agents for agentic tasks.
@@ 6 Discussion, Limitations & Conclusion
The ability for LLMs to leverage test-time compute has been transformative, yet efficiently allocating these resources in agentic settings stands as a critical, largely unexplored challenge. To the best of our knowledge, this work presents the first systematic investigation and learned solution enabling LLM agents to effectively allocate test-time compute in sequential decision-making tasks. Our findings reveal that prevailing always-plan (e.g. ReAct) and never-plan approaches are suboptimal. Instead, a "Goldilocks" effect emerges whereby intermediate frequencies outperform both extremes, likely avoiding instability ($C_{noise}$ in our framework) induced by excessive replanning. This highlights the need for LLM agents to strategically allocate, rather than naively scale, deliberation resources. Our two-stage SFT+RL methodology demonstrates that agents can learn this meta-cognitive skill, moving beyond fixed heuristics towards the adaptive, efficient behaviour essential for sustained autonomy. Moreover, the resulting agents become sufficiently adept at planning and execution to be effectively steered by human-written plans towards remarkable feats, including the full completion of Crafter through diamond collection, significantly impacting human-AI collaboration and safety.
@@
In summary, this paper establishes that in zero-shot evaluation, per-timestep planning or reasoning strategies akin to ReAct are outperformed by “Goldilocks” planning frequencies, likely due to instability that results from excessive planning or “overthinking.” By using a two-stage training methodology combining SFT and RL, we successfully train agents to dynamically allocate planning resources at test-time. This approach yields more effective and efficient behaviour compared to fixed planning strategies, marking a step towards more autonomous and scalable agentic systems.

### v2

@@ 1 Introduction
Figure 1: Dynamic planning strategies across environments and training stages. (a-b) Zero-shot results showing optimal “Goldilocks” planning frequency in Crafter and POGS (100 seeds, bars=standard-error). (c-d) SFT results demonstrating planning agents’ improved performance with lower KL divergence from base model. (e-f) RL results where SFT-primed planning agents are more sample efficient than non-planning baselines and more consistently reach complex achievements.
@@
    Each task has a “Goldilocks” frequency for planning that clearly outperforms naive strategies of always planning or never planning.
@@
Our work provides clear evidence that dynamic planning facilitates effective allocation of test-time compute in sequential decision making environments, showing that LLM agents can be trained to use additional computational resources intelligently. The ability to steer such planning agents—now capable enough to complete Crafter by collecting diamonds under human guidance—marks a significant step towards safer and more collaborative LLM agents. Together, these findings suggest a promising path towards more capable, efficient, interpretable, and steerable agentic systems.
@@ 2 Related Work
Historically, much progress in sequential decision making has involved systems that explicitly look ahead before acting. Monte Carlo Tree Search (MCTS)(Coulom, 2006), combined with deep neural networks, has driven landmark systems like AlphaGo and MuZero(Silver et al., 2017b; a; Schrittwieser et al., 2020). Model Predictive Control (MPC) iteratively plans short horizons, adapting to new observations (Mayne et al., 2000). Similarly, model-based RL methods, such as World Models, PlaNet, and Dreamer, use imagination rollouts in latent space for effective planning and learning (Ha & Schmidhuber, 2018; Hafner et al., 2019; 2020). Collectively, these approaches underscore the strength of explicit planning, particularly when accurate internal or environmental models are available.
@@
Large language models have demonstrated significant reasoning capabilities, particularly through techniques like chain-of-thought (CoT) prompting (Wei et al., 2022). ReAct extends CoT into sequential settings, explicitly prompting models to reason before acting (Yao et al., 2023b). Similar reasoning methods include self-reflective prompting (Shinn et al., 2023; Wang et al., 2023; Hao et al., 2023; Besta et al., 2024; Yao et al., 2023a), automated prompt tuning (Fernando et al., 2024; Hu et al., 2025), and strategic planning demonstrated by CICERO in the Diplomacy game († et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu, et al., FAIR). However, frequent replanning can cause behavioural instability, analogous to RL frame-skipping strategies that advocate less frequent action repetition for improved exploration and consistency (Sharma et al., 2017; Kalyanakrishnan et al., 2021). Recent studies also show diminishing returns and increased brittleness from excessive reasoning (Stechly et al., 2024; Mizrahi et al., 2024; Sui et al., 2025), emphasizing the need for adaptive planning mechanisms. Behaviour cloning with LLMs on data that includes textual reasoning between actions has been shown to help imitation learning (Yang et al., 2022; Hu & Clune, 2023).
@@ 4.1 Environments
To evaluate dynamic planning across different conditions, we select two complementary environments. First, Partially Observable Graph Search (POGS) is our custom synthetic environment designed to isolate planning under uncertainty. Agents navigate procedurally generated graph using only local observations, which require adaptive replanning upon discovery of new nodes or dead ends. POGS allows measurement of exploration efficiency via backtracking statistics. Second, Crafter (Hafner, 2022) is a complex 2D grid-world, long-horizon benchmark inspired by Minecraft. It demands multi-scale planning for survival, resource management, and crafting, testing both short-term tactical decisions and long-term strategic choices. Interaction in both environments occurs via natural language. Full technical details and figures for the environments are provided in the Appendix B.
@@ 4.2 Evaluation Protocol
We utilize the BALROG benchmark (Paglieri et al., 2025a) for standardized agent evaluation and environment interaction. At each timestep $t$, the agent receives its history and current observation $o_{t}$ within a chat template, guided by a system prompt outlining the task (Fig. 4 in Appendix A). The agent’s response must include a natural language action command $a_{t}$. Our dynamic planning agents are instructed to decide at each step whether to plan. If they choose not to plan, they output only the action command [Action]. If they choose to plan, they output the plan followed by the action, using the format <plan> [natural language plan] </plan> [Action]. BALROG parses this output, identifying a planning decision ($d_{t}=1$) if the <plan> block is present and using its content as the current plan $p_{t}$ in subsequent context. The [Action] command is always extracted and executed. Fallback mechanisms ensure robustness against invalid outputs. Appendix A provides detailed prompts.
@@ 4.4 Supervised Fine-Tuning (SFT)
SFT Priming: The Llama-3.1-8B model was fine-tuned on this data, aligning the SFT process with the target RL configuration. For the SFT+RL plan dynamically agent, SFT targets use the dynamic format – <plan>...</plan> [Action] if the teacher planned at that step, otherwise just [Action] – along with a dynamic prompt encouraging this choice. Conversely, for the SFT+RL no plan agent, SFT targets include only the actions ([Action], with all plan blocks removed), paired with a naive prompt focused solely on action prediction. This prepares the model appropriately for the subsequent RL phase. More details on the prompt in the Appendix A.
@@ 4.5 Reinforcement Learning (RL)
We then used Proximal Policy Optimization (PPO) (Schulman et al., 2017) to fine-tune Llama-3.1-8B-Instruct agents in Crafter, optimizing task rewards possibly adjusted for planning costs (Sec. 3.3). We compare four key configurations:
@@ 5.1 Zero-Shot Evaluation
To establish baseline capabilities and characterize the trade-offs between planning frequency, computational cost, and task performance, we conduct zero-shot evaluations. Specifically, we assess the zero-shot performance of Llama-3.3-70B-Instruct in the POGS and Crafter environments over 100 seeds. We systematically vary the agent’s planning frequency, from never planning to planning at fixed intervals, and measured task progression against the mean number of output tokens (log scale), which serves as a proxy for the computational budget.
Contrary to the intuitive assumption that performance would scale monotonically with computational effort to form a Pareto frontier, our findings reveal a non-monotonic relationship, which we term the “Goldilocks” zone for planning frequency. As shown in Figure 1 (a-b), performance in both environments peaks at intermediate planning frequencies before declining as planning becomes more frequent. The degradation in performance associated with excessive planning aligns with the instability cost ($C_{noise}$) introduced in our conceptual framework (Section 3). Quantitative analysis of agent trajectories in the POGS environment (Appendix B) supports this view; the always-plan agent exhibited the highest rate of backtracking, suggesting a tendency to oscillate or revisit states rather than explore efficiently. We hypothesize that this phenomenon may be compounded by other factors. For instance, excessive reasoning can lead to “overthinking” Sui et al. (2025) and longer contexts, which in turn dilute attention to the most critical information Liu et al. (2023).
Finally, it is noteworthy that a dynamic planning baseline is absent from our zero-shot analysis. Our initial attempts to elicit adaptive planning directly through complex prompting proved challenging and unreliable. Models struggled to consistently interpret abstract instructions such as “plan only when necessary,” often defaulting to fixed patterns of always planning or never planning. This difficulty underscores the need for learning-based approaches, like SFT and RL, to effectively teach agents this meta-cognitive skill of deciding when to allocate test-time compute for planning.
@@ 5.2 SFT Priming
Having established in zero-shot evaluations that prompting alone fails to induce dynamic planning, we next turn to supervised fine-tuning. Recent work shows that RL tends to optimize within the existing behaviors of the model rather than induce new strategies Ma et al. (2025); Gandhi et al. (2025). This limitation underscores the necessity of an SFT priming stage that provides explicit demonstrations over a mixture of both planning and non-planning steps within a trajectory.
To evaluate the impact of SFT priming, we leverage the synthetic dataset introduced in Section 4.4. During this phase the models were trained either to predict both the explicit plans and corresponding actions (‘Primed-Dynamic’) or to predict only the actions (‘Primed-Naive’). Importantly, both variants were derived from the same underlying action sequences, isolating the effect of explicit plans on imitation learning.
Figure 1c–d shows that model fine-tuned on the dynamic planning dataset (“Primed-Dynamic”) achieve significantly higher task progression throughout training compared to model trained without plans (“Primed-Naive”). Moreover, the Primed-Dynamic agent exhibits lower KL divergence from the base model, suggesting that the inclusion of plans regularizes the fine-tuning process and prevents catastrophic forgetting. We hypothesize three complementary mechanisms underlying these improvements:
@@
  • Explanatory Power: Plans provide semantic rationales for actions, simplifying the behavioral cloning objective and making future actions more predictable.
  • Learning Planning Structure: Exposure to explicit plan–action pairs encourages the model to internalize generalizable planning heuristics rather than memorizing isolated actions.
@@
  • Regularization and Grounding: Natural language plans act as a form of grounding signal, constraining the model’s learning trajectory and reducing deviation from base capabilities.
Collectively, these findings indicate that explicit plan representations not only improve downstream RL sample efficiency but also enhance the SFT stage itself by stabilizing learning dynamics and accelerating task progression.
5.3 RL Fine-Tuning and Instruction Following
Following the supervised fine-tuning stage, we apply RL to further refine the agents’ capabilities. The results (Fig. 1e-f and Fig. 3) from this phase highlight two main findings: (i) SFT priming is essential prerequisite for learning effective dynamic planning under RL; and (ii) the two-stage training process yields agents that can be steered by human-written plans to accomplish tasks beyond their autonomous performance.
Our experiments reveal that the SFT priming stage is crucial for enabling agents to develop dynamic planning skills through RL. The evidence for this is not just that SFT-primed agents outperform their non-primed counterparts, but is most starkly demonstrated by the fact that the Base+RL plan dynamically agent performed worse than the Base+RL no plan agent. This suggests that without the structured demonstrations provided by SFT, the base model cannot learn to plan effectively through RL alone; the attempt to generate plans is not only unhelpful but actively degrades performance. SFT priming is what provides the essential scaffolding, turning planning from a detrimental behavior into a highly effective strategy.
The agent primed with explicit plans (SFT+RL plan dynamically) significantly outperforms the agent primed only on actions (SFT+RL no plan). When planning is layered on top of SFT foundation, we observe improved sample efficiency and task progression. Qualitative analysis confirms that this agent successfully learns to generate and execute plans at varying levels of abstraction, demonstrating the ability to plan when necessary and replan in response to changing circumstances. We further observe that the planning cost penalty $C_{tokens}$ introduced in Section 3.3 effectively adjusts agent behavior in proportion to the penalty magnitude, with higher costs leading to reduced planning frequency and length during training while preserving task performance, indicating that agents can flexibly adapt their planning strategies to different computational constraints (see Appendix C.3 for detailed results).
Steerability and Human-Agent Collaboration While the training methodology produces significant performance improvements, computational constraints prevent any of the autonomous agents from fully solving the Crafter environment. To better understand the potential upper bound of our approach, we also evaluate agents in a human-in-the-loop setup.
The SFT+RL plan dynamically agent proves to be the most reliable at instruction following, making fewer mistakes during combat and adhering more consistently to the plans than the base and SFT-only models. Under human‑provided plans, it successfully completes Crafter by collecting diamonds—an achievement not observed in autonomous runs (see Fig. 3 for a representative sequence). Additional qualitative examples and a Best‑of‑N analysis are provided in the Appendix D.
@@ 6 Discussion, Limitations & Conclusion
The ability for LLMs to leverage test-time compute has been transformative, yet efficiently allocating these resources in agentic settings stands as a critical, largely unexplored challenge. To the best of our knowledge, this work presents the first systematic investigation and learned solution enabling LLM agents to effectively allocate test-time compute in sequential decision-making tasks. Our findings reveal that prevailing always-plan (e.g. ReAct) and never-plan approaches are suboptimal. Instead, a “Goldilocks” effect emerges whereby intermediate frequencies outperform both extremes, likely avoiding instability ($C_{noise}$ in our framework) induced by excessive replanning. This highlights the need for LLM agents to strategically allocate, rather than naively scale, deliberation resources. Our two-stage SFT+RL methodology demonstrates that agents can learn this meta-cognitive skill, moving beyond fixed heuristics towards the adaptive, efficient behaviour essential for sustained autonomy. Moreover, the resulting agents become sufficiently adept at planning and execution to be effectively steered by human-written plans towards remarkable feats, including the full completion of Crafter through diamond collection, significantly impacting human-AI collaboration and safety.
@@
In summary, this paper establishes that in zero-shot evaluation, per-timestep planning or reasoning strategies akin to ReAct are outperformed by “Goldilocks” planning frequencies, due to instability that results from excessive planning or “overthinking.” By using a two-stage training methodology combining SFT and RL, we successfully train agents to dynamically allocate planning resources at test-time. This approach yields more effective and efficient behaviour compared to fixed planning strategies, marking a step towards more autonomous and scalable agentic systems.
Ethics Statement
This work studies dynamic planning in language model agents within simulated environments, without human subjects or personal data. While improved planning could be misused in real-world autonomous systems, our evaluations are confined to sandboxed benchmarks, and we believe the scientific benefits outweigh these risks.
Reproducibility Statement
We strive to make all experiments in this paper fully reproducible. We share the anonymous codebase at: https://anonymous.4open.science/r/LLM-Agents-Planning. We provide further descriptions of the training details in Appendix C.